OpenStudy (anonymous):

Solve the differential equation dy dt = t(y2 − 4) by separation of variables. The integrals involved may require the use of partial fractions.

OpenStudy (anonymous):

do you mean dy/dt=t(y^2-4)?? sorry i cant understand

OpenStudy (anonymous):

ya sorry

OpenStudy (anonymous):

\[\frac{dy}{dt}=t(y^2-4) \implies \frac{dy}{y^2-4}=tdt \implies \int\limits\frac{dy}{y^2-4}=\int\limits tdt.\] Evaluate the integrals and don't forget to add the constant.

OpenStudy (anonymous):

the end result would be \[y=\pm \sqrt{(Ae^t^2)+4}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

When evaluating the integrals you will should get \[\frac{1}{4}\ln {2-y \over 2+y}=\frac{1}{2}t^2+c \implies \ln {2-y \over 2+y}=2t^2+c \implies {2-y \over 2+y}=ke^{2t^2}\] Solve for y now, if you want.

OpenStudy (anonymous):

It can also be written as \[\frac{2-y}{2+y}=Ae^{t^2}\] where \(A=ke^2\).

OpenStudy (anonymous):

i dont understand how you get this part 1/4 ln(2−y/2+y)

OpenStudy (anonymous):

By evaluating the Left-hand-side integral (\(\large \int {dy \over y^2-4}\)), where you need to use partial fractions.\[\]